Find the sample variance for the following data set 23 14 32 29 17 18 58.5 076 46.8
Answer #1
X x-x¯ (x-x¯)²
23 0 0
14 -9 81
32 9 81
29 6 36
17 -6 36
Total 115 ∑(x-x¯)²=234
Mean(x¯) 115/5=23

\\ Variance=\frac{\sum (x-\bar{x})^2}{n-1}=\frac{234}{4}=58.5

Ans: 58.5